solve the equation by factoring. z^2-10z+9=0

Answer by stanbon(75887)

solve the equation by factoring.
z^2-10z+9=0
----
z^2 -9z-z + 9 = 0
-----
z(z-9)-(z-9) = 0
-----
(z-9)(z-1) = 0
-----
z = 9 or z = 1
====================
Cheers,
Stan H.
==================